1.evaluate 5x-2y +4z when x=3, y=2 and z=4.
2.evaluate 6a+4b-3c when a=4, b=7, and c=-2.
3.evaluate 3x+5y when x=11 and y=4.

Answer by HyperBrain(694)

1.evaluate 5x-2y +4z when x=3, y=2 and z=4.
We will just substitute x=3, y=2 and z=4
So,

5x-2y+4z=5(3)-2(2)+4(4)
________=15-4+16
________=(15-4)+16
________=11+16
________=27


2.evaluate 6a+4b-3c when a=4, b=7, and c=-2.
We will just substitute a=4, b=7, and c=-2.
So,
6a+4b-3c=6(4)+4(7)-3(-2)
________=24+28-(-6)
________=(24+28)-(-6)
________=52-(-6)
________=52+6
________=58

3.evaluate 3x+5y when x=11 and y=4.
We will simply substitute x=11 and y=4
So,
3x+5y=3(11)+5(4)
_____=33+20
_____=53


Power up,
HyperBrain!